A 20.0mL sample of concentrated sodium hydroxide (NaOH) solution was diluted to a final volume of 60.0mL and 4.75mol/L. What mass of solid sodium hydroxide was needed to make the concentrated solution?
1
Expert's answer
2020-12-08T05:34:24-0500

There are 2 solutions.

The concentrated solution and the diluted solution.

The equation connecting them is;

C1V1 = C2V2

with the subscript 1 for the concentrated solution and the subscript 2 for the diluted solution.


C1×20 = 4.75 × 60

C1 = 285/20

C1 = 14.25mol/L


number of moles of the concentrated solution = C1 × V1 = 14.25 × 20/1000 = 280/1000 = 0.285 mol.


Mass of a substance = molar mass × number of moles = 58.5 × 0.285 = 16.273g



Therefore, 16.273g of solid sodium hydroxide was needed to make the concentrated solution.